OpenStudy (anonymous):

Help please! Solve the equation for y: 4x + 3y = -12. A. y = -12 - 4/3 x B. y = -4 - 4/3 x C. y = -12 + 4x D. y = -4 + 4x

OpenStudy (anonymous):

its A

OpenStudy (anonymous):

4x + 3y = -12 3y=-12-4x y=(-12-4x)/3 y=-4-(4/3)x

OpenStudy (anonymous):

sorry!! B

OpenStudy (anonymous):

Thank you very much!